# Optimal Oblivious Load-Balancing for Sparse Traffic in Large-Scale Satellite Networks

### v1

@@ I Introduction
Oblivious routing policies are predetermined routes from sources to destinations that are independent of the underlying traffic, which aim to minimize the maximum load on any link. In this work, we are primarily motivated by the problem of oblivious load-balancing of traffic in large-scale Low-Earth Orbit (LEO) satellite constellations designed for data communication. We consider a satellite network in which each satellite maintains four intersatellite links (ISLs): two to neighboring satellites in the same orbit and two to satellites in adjacent orbits. These ISLs are relatively stable and easy to maintain [handley_delay_2018]. A common dynamic model for such networks is the snapshot model [werner_dynamic_1997, ekici_distributed_2001]. For simplicity, we assume that at each snapshot the network can be modeled as a toroidal mesh [sun_capacity_2003, sun_routing_2004, Ramakanth2025] (see Fig. 1). This assumption is reasonable because the ISL topology remains approximately constant over routing timescales [handley_delay_2018]. While the satellite layer is relatively regular and static, RF links to ground nodes are highly dynamic [handley_delay_2018]. We therefore abstract the RF layer by assuming that traffic enters and exits directly at satellite nodes, and focus exclusively on load balancing within the satellite layer. Routing in the RF layer is handled independently by each satellite and is beyond the scope of this work.
@@
Our objective is to develop oblivious routing strategies that effectively load-balance sparse traffic over the intersatellite network, which is an $N\times N$ torus network. Oblivious routing policies for structured networks have been widely researched in the past [applegate_making_2006, chitavisutthivong_optimal_2023]. For $N\times N$ torus networks, the optimal oblivious routing problem has been studied in [towles_throughput-centric_2003, dally_principles_2004, ramanujam_randomized_2013]. However, prior works mostly optimize for load over a class of traffic called the hose model[duffield_flexible_1999] wherein every node can produce a maximum of one unit of traffic destined to any other node, and every node can also receive a maximum of one unit of traffic from any other node. Under this class of traffic, the celebrated Valiant Load-Balancing[valiant_scheme_1982, zhang-shen_designing_2005] scheme applied to a $N\times N$ torus is known to be one of the simplest optimal oblivious routing strategies[dally_principles_2004, ramanujam_randomized_2013]. However, this is true only when all nodes in the network generate traffic.
@@
The rest of the paper is organized as follows. In Section II, we describe the model of the satellite network, routing policies and classes of traffic in detail, and formulate the oblivious maximum link load problem. Section III develops a weak cut-based lower bound for the maximum link load of any routing policy under $k$-sparse traffic, demonstrates the suboptimality of the Valiant Load-Balancing Scheme, and describes a sufficient condition on the optimal routing policy. The main results of the paper, the lower and upper bounds, are presented in Sections IV and V respectively. We discuss some numerical results and concluding remarks in Section VI. Proofs (and general results) in this paper are omitted to highlight the key ideas of this work while keeping space limitations in mind. Finally, although this work is focused primarily on satellite networks, the results can also be applied to torus networks commonly used in other domains such as interconnection networks, optical networks and datacenter networks.
@@ II-A Network Model and Notation
For any $u,v\in\{0,...,N-1\}\times\{0,...,N-1\}$, we define the “-” operation similarly. We also denote 0 to be the origin node wherever appropriate, and $v\triangleq 0-v$ for any $v\in V$. It is important to note that operators ++ and - form a closure over $V$, i.e., they form a map from $V\times V\to V$. Next, we define $e_{1}=(1,0)$ and $e_{2}=(0,1)$. Then, the set $A=\{e_{1},e_{2},-e_{1},-e_{2}\}$ denotes the set of directions along which we have links from any node $i\in V$. In other words, every node $i$ in the network has outgoing links to $\{i+e_{1},i+e_{2},i-e_{1},i-e_{2}\}$. Similarly, every node $i$ also has incoming links from $\{i+e_{1},i+e_{2},i-e_{1},i-e_{2}\}$. Therefore, we can denote every link in the network using the starting node $i$ and the direction $e\in A$ along which it is directed. More specifically, the set of links can be represented as $E=\{(i,i+e)\;|\;i\in V,e\in A\}$. Every directed edge in $E$ has capacity $c_{i,i+e}$ which we assume to be $1$ for all $(i,i+e)\in E$.
@@ II-C Maximum Load on any Link
If $\mathcal{D}$ is a convex polytope of traffic matrices, then one can pose the above optimization problem as a linear program[dally_principles_2004, towles_throughput-centric_2003, applegate_making_2006]. Therefore, for any convex polytope of traffic matrices $\mathcal{D}$, the above program can be efficiently solved for moderately sized networks. However, this can get computationally challenging as the network size increases. In the next subsection, we define a special class of traffic matrices that are sparse, which are of practical importance.
@@ II-D The Class of Sparse Traffic Matrices
In this work, we start with the hose model which is quite popular in literature[duffield_flexible_1999, kodialam_maximum_2006, towles_throughput-centric_2003]. It is a practical model for traffic, which assumes that any node in the network can produce or sink traffic only up to a finite maximum capacity. In our hose model, without loss of generality, we assume that each node can source or sink one unit of traffic, that is,
@@
The proof of Lemma 1 comes from showing that the vertices of the polytope of $\mathcal{D}_{k}$ are in $\mathcal{D}^{\prime}_{k}$, i.e., $\text{vertices}(\mathcal{D}_{k})\subset\mathcal{D}^{\prime}_{k}$. Using this fact and the principles of linear programming, it is easy to see that for any routing policy, there is a worst-case traffic in the set $\mathcal{D}_{k}$ that would also be the worst-case traffic in the set $\mathcal{D}^{\prime}_{k}$. Consequently, an oblivious routing policy is optimal over traffic class $\mathcal{D}_{k}$ if and only if it is an optimal oblivious routing policy over traffic class $\mathcal{D}^{\prime}_{k}$. The details are deferred to Appendix A. Hence, our mathematical development depends on, and and applies to, the convex class of $k$-limited traffic matrices $\mathcal{D}_{k}$, which also captures our $k$-sparse traffic of interest.
@@ II-E Problem Formulation
Observe that constraint (9a) requires finding a maximum over the set $\mathcal{D}_{k}$ for every feasible routing policy. This means that we have to solve a linear program every time we need to ensure that constraint (9a) is satisfied. However, this can be easily circumvented using strong duality-based techniques [applegate_making_2006, towles_throughput-centric_2003, chitavisutthivong_optimal_2023, kodialam_maximum_2006]. These works employ strong duality to repose the oblivious max-load problem as a linear program. Although the problem of finding the optimal routing is a linear program, the number of constraints and variables becomes prohibitively large to computationally solve even for a moderately sized network. To that end, we aim to bound the performance of oblivious routing for sparse traffic by exploiting the structure in the problem. In the next section, we develop a preliminary cut-based lower bound for the maximum load. Furthermore, we reduce the search space for optimal routing policies by invoking a sufficiency condition for optimality.
@@ III-A Cut-Based Lower Bound
From the result in [sun_capacity_2003] on minimum sized cuts in a torus, for $k\leq N^{2}/4$, the minimum cut of a vertex set of size $k$ is roughly $4\sqrt{k}$. This is because, the minimum cut set of size $k$ is roughly a square of side length $\sqrt{k}$[sun_capacity_2003]. This is visually depicted in Figure 3.
@@ III-B Suboptimality of the Valiant Load-Balancing Scheme
However, one can show that the VLB scheme is an optimal oblivious routing for $k$-sparse traffic when $k\geq N^{2}/2$ and achieves a maximum link load of $N/4$ for all traffic in $\mathcal{D}_{k}$, $k\geq N^{2}/2$. The details are provided in Appendix G.
@@ III-C Sufficiency Condition for Optimality
The $N\times N$ torus network exhibits high degree of symmetry. Intuitively, symmetry means that the network looks the same even after translation or reflection transformations. We formalize this notion of symmetry using automorphisms[chitavisutthivong_optimal_2023].
Definition 1 (Automorphism [chitavisutthivong_optimal_2023]).
@@
Lemma 2 (Adopted from [chitavisutthivong_optimal_2023, Theorem 1]).
@@
The proof of Lemma 2 comes from modifying [chitavisutthivong_optimal_2023, Theorem 1] for max load, described in Appendix B. The sufficient condition tells us that it is enough to search for the optimal routing within routing policies that satisfies certain symmetry constraints. This drastically reduces the search space as it provides more structure to the linear program that can be exploited.
@@
$\theta^{*}=\min_{g\in\mathcal{R}_{sym}}\max_{d\in\mathcal{D}_{k}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}},$ (11)
@@
Lemma 3 exploits the structure of the torus to give us the max load problem as a bilinear optimization problem. From von Neumann’s minimax equivalence theorem [v_neumann_zur_1928] for bilinear objectives (or equivalently, the strong duality of linear programs), we also have the following equivalence.
$\theta^{*}$ $=\max_{d\in\mathcal{D}_{k}}\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}.$ (12)
@@ IV Worst-Case Load Lower Bound
$\frac{\sqrt{2k}}{4}$ $\lesssim\max_{d\in\mathcal{D}_{k}}\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}=\theta^{*}.$
@@ V-A Local Load-Balancing Policy (LLB)
Recall that it is sufficient for us to describe a set of routes from the origin 0 to every node $t=(t_{x},t_{y})$ such that $t_{x}\leq t_{y}\leq N/2$. Then, from the consequences of automorphism invariance, the route from any source $s$ to any destination $s+t$ can be obtained from the transformation of the aforementioned set of routes. The LLB scheme with $r=2$ is visualized in Figure 7 for a $N\times N$ torus with $N=7$.
@@
(b) Second phase - Route traffic from $S_{r}(0)$ to $S_{r}(t)$ along edge-disjoint paths with two edge-disjoint paths emanating from each node in $S_{r}(0)$ and terminating at each node in $S_{t}(t)$.
@@
Lemma 4 comes from a special case of the celebrated Max-Flow Min-Cut theorem (also known as Menger’s Theorem[lawler_combinatorial_1976]) to count the maximum number of edge-disjoint paths from a set of nodes to another set of nodes. The proof involves observing that the cut size of $S_{r}(i)$ is always $8r+4$ for $N\leq 2$, and hence there must exist at least $8r$ paths from $S_{r}(0)$ to $S_{r}(t)$. Moreover, these paths can be easily found using multiple runs of breadth-first search, removing edges once a path from $S_{r}(0)$ to $S_{r}(t)$ is found successively. We defer the details to Appendix F.
V-B Modifications to LLB policy under specific conditions
As discussed in the previous subsection it suffices to describe the routing policy when $t_{x}\leq t_{y}\leq N/2$ as for other destinations, the routing can be obtained from automorphism transformations. In this section, we limit ourselves to $t_{x}\leq t_{y}\leq N/2$ for ease of presentation.
V-B1 Type-I Overlap
V-C Maximum Load under LLB Policy
@@
$\theta^{*}$ $=\min_{g\in\mathcal{R}_{sym}}\max_{d\in\mathcal{D}_{k}}\sum_{s,t}d_{s,s+t}g^{t}_{s,s+e_{1}}\lesssim\frac{\sqrt{2k}}{4},$
@@
because the LLB scheme is a feasible routing, i.e., $g^{LLB}\in\mathcal{R}_{sym}$. Therefore, Theorems 1 and 2 tell us that the LLB scheme is an optimal oblivious routing scheme for sparse traffic when $2k$ is a perfect square. The proof part (a) of Theorem 2 involves looking at the dual of the inner maximization problem of the minmax problem in (11) and exploiting its structure to characterize a “good” candidate routing policy. Subsequently, we show that LLB is in fact a a “good” candidate routing policy which gives us a the required upper bound. The details are omitted here in the interest of space and can be found in Appendix E.
VI Numerical Results and Conclusion

### v2

@@ I Introduction
Oblivious routing policies are predetermined routes from sources to destinations that are independent of the underlying traffic, which aim to minimize the maximum load on any link. In this work, we are primarily motivated by the problem of oblivious load-balancing of traffic in large-scale Low-Earth Orbit (LEO) satellite constellations designed for data communication. We consider a satellite network in which each satellite maintains four intersatellite links (ISLs): two to neighboring satellites in the same orbit and two to satellites in adjacent orbits. These ISLs are relatively stable and easy to maintain [8]. A common dynamic model for such networks is the snapshot model [21, 7]. For simplicity, we assume that at each snapshot the network can be modeled as a toroidal mesh [17, 16, 14] (see Fig. 1). This assumption is reasonable because the ISL topology remains approximately constant over routing timescales [8]. While the satellite layer is relatively regular and static, RF links to ground nodes are highly dynamic [8]. We therefore abstract the RF layer by assuming that traffic enters and exits directly at satellite nodes, and focus exclusively on load balancing within the satellite layer. Routing in the RF layer is handled independently by each satellite and is beyond the scope of this work.
@@
While there have been multiple works studying load-balancing in satellite networks [9, 7, 17, 13, 12, 5, 14], they mostly do so with restrictive assumptions on traffic or using queue-based estimates of traffic. Our focus is on developing predetermined routing strategies that guarantee optimal worst-case load for sparse traffic, with only minimal assumptions on the underlying traffic pattern over the intersatellite network (which we model as an $N\times N$ torus network). Oblivious routing policies for structured networks have been widely researched in the past [1, 3]. For $N\times N$ torus networks, the optimal oblivious routing problem has been studied in [18, 4, 15]. However, prior works mostly optimize for load over a class of traffic called the hose model[6] wherein every node can produce a maximum of one unit of traffic destined to any other node, and every node can also receive a maximum of one unit of traffic from any other node. Under this class of traffic, the celebrated Valiant Load-Balancing[20, 22] scheme applied to a $N\times N$ torus is known to be one of the simplest optimal oblivious routing strategies[4, 15]. However, this is true only when all nodes in the network generate traffic.
@@
The rest of the paper is organized as follows. In Section II, we describe the model of the satellite network, routing policies and classes of traffic in detail, and formulate the oblivious maximum link load problem. Section III develops a weak cut-based lower bound for the maximum link load of any routing policy under $k$-sparse traffic, demonstrates the suboptimality of the Valiant Load-Balancing Scheme, and describes a sufficient condition on the optimal routing policy. The main results of the paper, the lower and upper bounds, are presented in Sections IV and V respectively. We discuss some numerical results and concluding remarks in Section VIII. Proofs (and general results) in this paper are omitted to highlight the key ideas of this work while keeping space limitations in mind. Finally, although this work is focused primarily on satellite networks, the results can also be applied to torus networks commonly used in other domains such as interconnection networks, optical networks and datacenter networks.
@@ II-A Network Model and Notation
For any $u,v\in\{0,...,N-1\}\times\{0,...,N-1\}$, we define the “-” operation similarly. We also denote 0 to be the origin node wherever appropriate, and $v\triangleq 0-v$ for any $v\in V$. It is important to note that operators ++ and - form a closure over $V$, i.e., they form a map from $V\times V\to V$. Next, we define $e_{1}=(1,0)$ and $e_{2}=(0,1)$. Then, the set $A=\{e_{1},e_{2},-e_{1},-e_{2}\}$ denotes the set of directions along which we have links from any node $i\in V$. In other words, every node $i$ in the network has outgoing links to $\{i+e_{1},i+e_{2},i-e_{1},i-e_{2}\}$. Similarly, every node $i$ also has incoming links from $\{i+e_{1},i+e_{2},i-e_{1},i-e_{2}\}$. Therefore, we can denote every link in the network using the starting node $i$ and the direction $e\in A$ along which it is directed. More specifically, the set of links can be represented as $E=\{(i,i+e)\;|\;i\in V,e\in A\}$. Every directed edge in $E$ has capacity $c_{i,i+e}$ which we assume to be $1$ for all $(i,i+e)\in E$. Finally, we use the terms “link” and “edge” interchangeably to refer to $(i,i+e)\in E$.
@@ II-C Maximum Load on any Link
If $\mathcal{D}$ is a convex polytope of traffic matrices, then one can pose the above optimization problem as a linear program[4, 18, 1]. Therefore, for any convex polytope of traffic matrices $\mathcal{D}$, the above program can be efficiently solved for moderately sized networks. However, this can get computationally challenging as the network size increases. In the next subsection, we define a special class of traffic matrices that are sparse, which are of practical importance.
@@ II-D The Class of Sparse Traffic Matrices
In this work, we start with the hose model which is quite popular in literature[6, 10, 18]. It is a practical model for traffic, which assumes that any node in the network can produce or sink traffic only up to a finite maximum capacity. In our hose model, without loss of generality, we assume that each node can source or sink one unit of traffic, that is,
@@
The proof of Lemma 1 comes from showing that the vertices of the polytope of $\mathcal{D}_{k}$ are in $\mathcal{D}^{\prime}_{k}$, i.e., $\text{vertices}(\mathcal{D}_{k})\subset\mathcal{D}^{\prime}_{k}$. Using this fact and the principles of linear programming, it is easy to see that for any routing policy, there is a worst-case traffic in the set $\mathcal{D}_{k}$ that would also be the worst-case traffic in the set $\mathcal{D}^{\prime}_{k}$. Consequently, an oblivious routing policy is optimal over traffic class $\mathcal{D}_{k}$ if and only if it is an optimal oblivious routing policy over traffic class $\mathcal{D}^{\prime}_{k}$. The details are deferred to Appendix A. Hence, our mathematical development depends on, and applies to, the convex class of $k$-limited traffic matrices $\mathcal{D}_{k}$, which also captures our $k$-sparse traffic of interest.
@@ II-E Problem Formulation
Observe that constraint (9a) requires finding a maximum over the set $\mathcal{D}_{k}$ for every feasible routing policy. This means that we have to solve a linear program every time we need to ensure that constraint (9a) is satisfied. However, this can be easily circumvented using strong duality-based techniques [1, 18, 3, 10]. These works employ strong duality to repose the oblivious max-load problem as a linear program. Although the problem of finding the optimal routing is a linear program, the number of constraints and variables becomes prohibitively large to computationally solve even for a moderately sized network. To that end, we aim to bound the performance of oblivious routing for sparse traffic by exploiting the structure in the problem. In the next section, we develop a preliminary cut-based lower bound for the maximum load. Furthermore, we reduce the search space for optimal routing policies by invoking a sufficiency condition for optimality.
@@ III-A Cut-Based Lower Bound
From the result in [17] on minimum sized cuts in a torus, for $k\leq N^{2}/4$, the minimum cut of a vertex set of size $k$ is roughly $4\sqrt{k}$. This is because, the minimum cut set of size $k$ is roughly a square of side length $\sqrt{k}$[17]. This is visually depicted in Figure 3.
@@ III-B Suboptimality of the Valiant Load-Balancing Scheme
However, one can show that the VLB scheme is an optimal oblivious routing for $k$-sparse traffic when $k\geq N^{2}/2$ and achieves a maximum link load of $N/4$ for all traffic in $\mathcal{D}_{k}$, $k\geq N^{2}/2$. The details are provided in Appendix H.
@@ III-C Sufficiency Condition for Optimality
The $N\times N$ torus network exhibits high degree of symmetry. Intuitively, symmetry means that the network looks the same even after translation or reflection transformations. We formalize this notion of symmetry using automorphisms[3].
Definition 1 (Automorphism [3]).
@@
Lemma 2 (Adopted from [3, Theorem 1]).
@@
The proof of Lemma 2 comes from modifying [3, Theorem 1] for max load, described in Appendix B. The sufficient condition tells us that it is enough to search for the optimal routing within routing policies that satisfies certain symmetry constraints. This drastically reduces the search space as it provides more structure to the linear program that can be exploited.
@@
$\theta^{*}=\min_{g\in\mathcal{R}_{sym}}\max_{d\in\mathcal{D}_{k}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}},$ (11)
@@
Lemma 3 exploits the structure of the torus to give us the max load problem as a bilinear optimization problem. From von Neumann’s minimax equivalence theorem [19] for bilinear objectives (or equivalently, the strong duality of linear programs), we also have the following equivalence.
$\theta^{*}$ $=\max_{d\in\mathcal{D}_{k}}\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}}.$ (12)
@@ IV Worst-Case Load Lower Bound
$\frac{\sqrt{2k}}{4}$ $\lesssim\max_{d\in\mathcal{D}_{k}}\min_{g\in\mathcal{R}_{sym}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}}=\theta^{*}.$
@@ V Optimal Oblivious Routing for Sparse Traffic
Recall that it is sufficient for us to describe a set of routes from the origin 0 to every node $t=(t_{x},t_{y})$ such that $t_{x}\leq t_{y}\leq N/2$. Then, from the consequences of automorphism invariance, the route from any source $s$ to any destination $s+t$ can be obtained from the transformation of the aforementioned set of routes. When there is overlap between the stems $S_{r}(0)$ and $S_{r}(t)$, minor modifications need to be made in the first and third phase, which we omit in the interest of space. Refer to Appendix F for more detail. The LLB scheme with $r=2$ is visualized in Figure 7 for a $N\times N$ torus with $N=7$.
@@
(b) Second phase - Route traffic from $S_{r}(0)$ to $S_{r}(t)$ along edge-disjoint paths with two edge-disjoint paths emanating from each node in $S_{r}(0)$ and terminating at each node in $S_{r}(t)$.
@@
Lemma 4 comes from a special case of the celebrated Max-Flow Min-Cut theorem (also known as Menger’s Theorem[11]) to count the maximum number of edge-disjoint paths from a set of nodes to another set of nodes. The proof involves observing that the cut size of $S_{r}(i)$ is always $8r+4$ for $r<N/2$, and hence there must exist at least $8r$ paths from $S_{r}(0)$ to $S_{r}(t)$. Moreover, these paths can be easily found using multiple runs of breadth-first search, removing edges once a path from $S_{r}(0)$ to $S_{r}(t)$ is found successively. We defer the details to Appendix G.
@@
$\theta^{*}$ $=\min_{g\in\mathcal{R}_{sym}}\max_{d\in\mathcal{D}_{k}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}}\lesssim\frac{\sqrt{2k}}{4},$
@@
because the LLB scheme is a feasible routing, i.e., $g^{LLB}\in\mathcal{R}_{sym}$. Therefore, Theorems 1 and 2 tell us that the LLB scheme is an optimal oblivious routing scheme for sparse traffic when $2k$ is a perfect square. To prove Theorem 2, we first greedily maximize the load any link can experience under the LLB policy and show that it is bounded above by $\sqrt{2k}/4$ when $2k$ is a perfect square. Later, we use the result of part (a) to show part (b). The details are omitted here in the interest of space and can be found in Appendix E.
VI Extension to General $N\times M$ Tori
In this section, we extend our results to $N\times M$ tori with capacities $c_{1}$ along the horizontal links and $c_{2}$ along the vertical links, as shown in Figure 8. The $N\times M$ torus has links at each node along the directions in set $A=\{e_{1},-e_{1},e_{2},-e_{2}\}$, where $e_{1}=(0,1)$ is the positive vertical direction, $e_{1}=(0,-1)$ is the negative vertical direction, $e_{2}=(1,0)$ is the positive horizontal direction, and $e_{2}=(-1,0)$ is the negative horizontal direction. The links in the vertical directions, that is, along $e_{1}$ and $e_{1}$ have a capacity of $c_{1}$, and the links in the horizontal directions, that is, along $e_{2}$ and $e_{2}$ have a capacity of $c_{2}$. The links along $\pm e_{1}$ and $\pm e_{2}$ are not identical, unlike the case in the previous sections.
We are interested in developing the optimal oblivious load balancing strategy for the general asymmetric torus network topology. In other words, we want to characterize an optimal routing strategy $f^{*}$ such that
$\max_{d\in\mathcal{D}_{k}}\textsc{MaxLoad}(f^{*},d)=\min_{f\in\mathcal{R}}\max_{d\in\mathcal{D}_{k}}\textsc{MaxLoad}(f,d)$ (16)
for the general $N\times M$ torus.
Figure 8: $N\times M$ torus with $N=4$ and $M=5$.
Observe that $N\times M$ torus, exhibits translation symmetry and symmetry under reflection about the origin. However, unlike the $N\times N$ torus, the $N\times M$ torus does not exhibit symmetry under reflection about the $x=y$ axis. This is the only major difference between the analysis in the previous sections and in this section. Using the automorphisms of the general $N\times M$ torus, we can formulate a reduced version of problem (16), similar to that of Lemma 3 for a $N\times N$ torus. Like in Section III-C, we will again denote the reduced routing policy as $g^{t}_{i,i+e}$. Essentially, $g^{t}_{i,i+e}$ represents the fraction of traffic from the origin to node $t$ routed along the edge $(i,i+e)$. Using this characterization, one can obtain the fraction of any $(s,s+t)$ traffic routed along any edge $(i,i+e)$ simply by using the translation transformation $f^{s,s+t}_{i,i+e}=g^{t}_{i-s,i-s+e}$. It is sufficient to characterize an optimal route from the origin to every destination $t$ to describe an optimal route for all source-destination pairs. The reduced problem for the general $N\times M$ torus is stated in Lemma 5.
Lemma 5.
$\min_{g,\theta}\quad$ $\theta$ (17a)
s.t. $\max_{d\in\mathcal{D}_{k}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{1}}\leq c_{1}\theta$ (17b)
$\max_{d\in\mathcal{D}_{k}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{2}}\leq c_{2}\theta$ (17c)
$\begin{aligned} \sum_{e\in A}g^{t}_{i,i+e}-&\sum_{e\in A}g^{t}_{i-e,i}=\begin{cases}0&i\neq 0,t\\ 1&i=0\\ -1&i=t\end{cases}\end{aligned}$ $\forall i,t$ (17d)
$\begin{aligned} g^{t}_{i,i+e}&=g^{R_{0}(t)}_{R_{0}(i),R_{0}(i+e)}\end{aligned}$ $\forall i,e,t$ (17e)
$0\leq g^{t}_{i,i+e_{1}}\leq 1$ $\forall i,t.$ (17f)
Problem (17) can be concisely stated as
$\theta^{*}=\min_{g\in\mathcal{R}^{\prime}_{sym}}\max_{d\in\mathcal{D}_{k}}\max_{j\in\{1,2\}}\frac{1}{c_{j}}\sum_{s,t}d_{s,s+t}g^{t}_{-s,-s+e_{j}},$ (18)
where $\mathcal{R}^{\prime}_{sym}$ is the class of routing policies which satisfy translation and reflection invariance, i.e.,
[…]
VI-A Worst-Case Load Lower Bound for general $N\times M$ tori
We use insights developed from the proof of Theorem 1 to construct a lower bound for (17). For the general $N\times M$ torus, we consider a generalized version of the Split-Diamond traffic in order to develop lower bounds on the load that any routing policy in $\mathcal{R}^{\prime}_{sym}$ would have to incur. This is formally stated in Theorem 3.
Theorem 3.
For a $N\times M$ torus with capacities $c_{1}$ along the vertical links and $c_{2}$ along the horizontal links
a) When $k\leq L^{2}/2$, all routing policies $g\in\mathcal{R}^{\prime}_{sym}$ incur a maximum load of at least $\frac{\sqrt{2k}}{4\sqrt{c_{1}c_{2}}}-\epsilon^{\prime}$,
b) When $L^{2}/2\leq k\leq NM/2$, all routing policies $g\in\mathcal{R}^{\prime}_{sym}$ incur a maximum load of at least $\frac{k}{2L\sqrt{c_{1}c_{2}}}$,
c) When $k\geq NM/2$, all routing policies $g\in\mathcal{R}^{\prime}_{sym}$ incur a maximum load of at least $\frac{NM}{4L\sqrt{c_{1}c_{2}}}$,
where $L=\min(\sqrt{c_{2}/c_{1}}N,\sqrt{c_{1}/c_{2}}M)$.
Theorem 3 immediately gives us a lower bound on
$\min_{f\in\mathcal{R}}\max_{d\in\mathcal{D}_{k}}\textsc{MaxLoad}(f,d)\gtrsim\begin{cases}\frac{\sqrt{2k}}{4\sqrt{c_{1}c_{2}}}&k\leq\frac{L^{2}}{2}\\ \frac{k}{2L\sqrt{c_{1}c_{2}}}&\frac{L^{2}}{2}\leq k\leq\frac{NM}{2}\\ \frac{NM}{4L\sqrt{c_{1}c_{2}}}&k\geq\frac{NM}{2}\end{cases}.$
$L$ is to be interpreted as the normalized size of smallest dimension of the torus. It is in fact the minimum bisection bandwidth of the general $N\times M$ torus normalized by 2 times the geometric mean of capacities along each dimension, i.e., $2\sqrt{c_{1}c_{2}}$. We prove Theorem 3 in Appendix I.
VII Optimal Oblivious Routing for Sparse Traffic in General $N\times M$ torus
In this section, we propose a set of routing policies that achieve the lower bound on max load described in Theorem 4. We define a generalization of the LLB scheme in Section V which we call the Generalized Local Load Balancing (GLLB) scheme. The GLLB scheme is parametrized by $r_{1},r_{2}\in\mathbb{N}$ and $r_{1}\leq N/2,r_{2}\leq M/2$. We denote the routing policy as $g^{GLLB(r_{1},r_{2})}$. The routing policy is specifically designed to near-optimally load balance any $k$-sparse traffic effectively in the $N\times M$ network for an appropriate choice of $r_{1}$ and $r_{2}$ when $k\leq L^{2}/2$. The structure of the GLLB policy is inherently dependent on the geometry of the network. As a result, it must be formulated on a case-by-case basis, with its precise form determined by the parameters $r_{1},r_{2}$, the network size $N,M$, and link capacities $c_{1}$ and $c_{2}$. When $k\geq L^{2}/2$, we have a simple routing strategy that achieves the lower bound. For the detailed description of GLLB, refer to Appendix J.
Theorem 4.
For a $N\times M$ torus with capacities $c_{1}$ along the vertical links and $c_{2}$ along the horizontal links
a) When $k\leq L^{2}/2$, Generalized Local Load-Balancing $g^{GLLB(r_{1},r_{2})}$, with parameter $r_{1}=\lceil\sqrt{c_{1}k/2c_{2}}\rceil$ and $r_{2}=\lceil\sqrt{c_{2}k/2c_{1}}\rceil$, achieves a load of at most $\frac{\sqrt{2k}}{4\sqrt{c_{1}c_{2}}}+\epsilon^{\prime}$ for any traffic in $\mathcal{D}_{k}$,
b) When $k\geq L^{2}/2$, there exists a routing strategy that achieves a load of at most $\min\{\frac{k}{2L\sqrt{c_{1}c_{2}}},\frac{NM}{4L\sqrt{c_{1}c_{2}}}\}$. for any traffic in $\mathcal{D}_{k}$,
where $L=\min(\sqrt{c_{2}/c_{1}}N,\sqrt{c_{1}/c_{2}}M)$.
Theorem 4 immediately gives us an upper bound on
$\min_{f\in\mathcal{R}}\max_{d\in\mathcal{D}_{k}}\textsc{MaxLoad}(f,d)\lesssim\begin{cases}\frac{\sqrt{2k}}{4\sqrt{c_{1}c_{2}}}&k\leq\frac{L^{2}}{2}\\ \frac{k}{2L\sqrt{c_{1}c_{2}}}&\frac{L^{2}}{2}\leq k\leq\frac{NM}{2}\\ \frac{NM}{4L\sqrt{c_{1}c_{2}}}&k\geq\frac{NM}{2}\end{cases}.$
The proof of Theorem 4 can be found in Appendix K.
VIII Numerical Results and Conclusion